Grade 11Mechanics

a body is projected up a 45degree rough incline. if the coefficient of friction is 0.5, then the retardation of the block is

Rishi Manoj
8 Years agoGrade 11

1 Answer

Venkat
8 Years ago
a body is projected up a 45degree rough incline. if the coefficient of friction is 0.5, then the retardation of the block is
 
Net retarding force = \\mgsin45+\mu cos45
 
Retardation  = F/m

\\retardation = \frac{mgsin45+\mu (mg cos45)}{m} \\ \\retardation =\frac{g}{\sqrt{2}}(1+0.5) \\ \\retardation =\frac{3g}{2\sqrt{2}}
                = (g/√2)*[1+0.5]
               = {3/(2√2)}*g